A tuning fork produces 5 beats per sec when sounded with a sonometer wire of length 20cm. Number of beats herd per second remains the same even if the length of sonometer wire is changed to 21 cm. The frequency of tuning fork is

Grade:12

2 Answers

Arun
6 years ago
Dear student
let the frequecy be = n
then the possible frequency of sonometer wire are = n + 5, n –5
n + 5 = (1/40) * sqrt (T/m),
n –5 =  (1/42) * sqrt (T/m)
Hence
n + 5/ n –5 = 21/ 20
20 n + 100 = 21 n – 105
n = 205 Hz
Regards
Arun (askIITians forum expert)
Rishi Sharma
3 years ago
Dear Student,
Please find below the solution to your problem.

Let the frequecy be = n
then the possible frequency of sonometer wire are = n + 5, n –5
n + 5 = (1/40) * sqrt (T/m),
n –5 = (1/42) * sqrt (T/m)
Hence n + 5/ n –5 = 21/ 20
20 n + 100 = 21 n – 105
n = 205 Hz
Thanks and Regards